OpenStudy (anonymous):

Add. (- 3) + (- 6) + 7 + (- 8) =

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow -a + -b = -(a + b) }\) \(\Large \color{Black}{\Rightarrow a - b = -(b - a) }\) Two identities.

Parth (parthkohli):

A negative plus negative is negative.

OpenStudy (anonymous):

-10.

OpenStudy (anonymous):

-a + -b = -(a+b) and then a-b = -(b-a) two negatives make it a positive, then you'd have to add

OpenStudy (anonymous):

-10

Parth (parthkohli):

For example if you have \(6 + (-8)\) Then you'll get \(6 - 8\) further \(\Large \color{Black}{\Rightarrow -(8 - 6) = -2}\)

Parth (parthkohli):

Or do them step-by-step. Add all the positives first, then the negatives.

OpenStudy (anonymous):

That's the easiest way for me

OpenStudy (anonymous):

thnxs